OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. open parentheses x to the 3 fourths power close parentheses to the 3 sevenths power x to the 9 elevenths power x to the 9 over 28 power x to the 6 over 28 power x to the 33 over 28 power

OpenStudy (jackellyn):

\[(x ^{3/4})^{3/7}\]

OpenStudy (anonymous):

im honestly so confused

OpenStudy (apoorvk):

@jackellyn appreciate your helping, but you made a fundamental mistake here.. Actually, it's ---> \( a^m \times a^n = a^{m+n}\) But, \((a^m)^n = a^{mn}\) brings back some old memories? =)

OpenStudy (apoorvk):

@snapback_champ can you use the funda posted above to solve the problem?

OpenStudy (jackellyn):

you're right, sorry @snapback_champ and @apoorvk honest mistake

OpenStudy (apoorvk):

No worries, keep going ;)

OpenStudy (anonymous):

i figured it out!!! thanks @apoorvk & @jackellyn